OpenStudy (anonymous):

need help solve this equation. r over 12+7=73

OpenStudy (anonymous):

is it (r/12) +7 = 73 or r/(12+7)=73?

OpenStudy (anonymous):

r/(12+7)=73 r=1387

OpenStudy (anonymous):

yes its a fraction

OpenStudy (anonymous):

both have fractions in the equation. which one specifically though.

OpenStudy (anonymous):

r 12

OpenStudy (anonymous):

ok. so you subtract the seven from both sides : r/12=66 then multiply by the denominator : r=66*12 which equals some large number that is easily found.

OpenStudy (anonymous):

thank u